### v1

# GRAF: Graph Retrieval Augmented by Facts for Legal Question Answering

Abstract
Pre-trained Language Models (PLMs) have shown remarkable performances in recent years, setting a new paradigm for NLP research and industry. The legal domain has received some attention from the NLP community partly due to its textual nature. Some tasks from this domain are represented by question-answering (QA) tasks. This work explores the legal domain Multiple-Choice QA (MCQA) for a low-resource language. The contribution of this work is multi-fold. We first introduce JuRO, the first openly available Romanian legal MCQA dataset, comprising three different examinations and a number of 10,836 total questions. Along with this dataset, we introduce CROL, an organized corpus of laws that has a total of 93 distinct documents with their modifications from 763 time spans, that we leveraged in this work for Information Retrieval (IR) techniques. Moreover, we are the first to propose Law-RoG, a Knowledge Graph (KG) for the Romanian language, and this KG is derived from the aforementioned corpus. Lastly, we propose a novel approach for MCQA, Graph Retrieval Augmented by Facts (GRAF), which achieves competitive results with generally accepted SOTA methods and even exceeds them in most settings.
GRAF: Graph Retrieval Augmented by Facts for Legal Question Answering
Cristian-George Craciun1,3thanks: cristian.craciun@tum.de Răzvan-Alexandru Smădu3, Dumitru-Clementin Cercel3thanks: dumitru.cercel@upb.ro, Mihaela-Claudia Cercel2,4 1Technical University of Munich, Munich, Germany 2Paris 1 Panthéon-Sorbonne University, Paris, France 3National University of Science and Technology POLITEHNICA Bucharest, Faculty of Automatic Control and Computers, Bucharest, Romania 4University of Bucharest, Bucharest, Romania
1 Introduction
With the advent of the Transformer architecture (vaswani2017attention) and its parallelization capabilities, people have sought to scale the models in terms of parameters (kaplan2020scaling), unlocking new abilities for machine learning models to perceive natural language at a finer level. Since then, Large Language Models (LLMs) have been pre-trained on various tasks, achieving better performances on downstream tasks (gururangan2020don) than their non-trained counterparts.
QA represents a family of NLP downstream tasks explored in various works (zhong2020jec; rogers2023qa; singhal2023towards). Some QA tasks can be formulated as open-ended questions that require an elaborate answer, which may or may not include the rationale behind it. Other works focus on retrieving the answer from a given context, in which case the answer can either be ad litteram found in the provided context (zaib2021bert) and a model has to identify the location at which the answer occurs, or the model has to infer the correct answer based on an understanding or interpretation of the context at hand, task which bears the name Machine Reading Comprehension (MRC) (baradaran2022survey).
The aforementioned tasks have been explored for various domains and languages (askari2022expert; louis2021statutory; hoppe2021towards). This work focuses on MCQA for the Romanian legal domain. Legal QA represents a very promising area of research due to the insights it can provide in tackling various other problems, as well as, from a practical standpoint, by benefiting the average citizen of a nation with the possibility of finding the answer to a legal inquiry together with experts in the field who will achieve increased productivity.
Additionally, available natural language data are scarce for the Romanian language. The only MCQA work was performed in (dima2024roqllama), which proposed a Romanian medical MCQA dataset. Although the legal domain for the Romanian language has been tackled in some works in the past (masala2021jurbert; masala-etal-2024-improving), there is no open-source datasets available. This motivated us to introduce new comprehensive resources for the Romanian language.
To summarize, the contributions of this paper are the following:
  • We release JuRO [^1] https://github.com/craciuncg/JuRO, the first open-source legal dataset available for the MCQA Romanian.
  • We release CROLfootnotemark: , a structured corpus of Romanian law that can be utilized to query the law for answers.
  • We release Law-RoGfootnotemark: , the first KG for the Romanian language, a KG for the legal field.
  • We present a novel algorithm for MCQA called GRAFfootnotemark: .
  • We provide a comprehensive evaluation of our proposed MCQA dataset using SOTA methods and our proposed method.
2 Related Work
A natural follow-up for the QA domain represented the adaptation of this research branch to downstream fields that require specific knowledge, particularly examinations. Also, the plethora of freely and publicly available legal texts made legal question answering a convenient start for testing diverse engineering solutions and targeted algorithms, allowing for the development of proof-of-concept methods.
2.1 Legal QA
English Work. There have been initial traditional approaches to legal question answering for bar exam questions (john2017solving) using BiLSTM networks. The remarkable achievements of the Transformer models (vaswani2017attention) motivated people to use them in this domain as well. One work revolved around privacy-related law shankar2023privacyglue, using various sources for building their benchmark dataset, besides using several legal BERT (devlin2018bert) models for their experiments. Another work has focused on IR of legal paragraphs, designing a new approach using a multi-stage pipeline, incorporating the BM25 algorithm (robertson1976relevance) and BERT for semantic similarity.
Multilingual Work. Other languages also took the initiative, and linguistic research groups attempted to solve legal tasks in their languages using diverse approaches. One work in the Chinese language addressed the legal field, featuring an MCQA dataset along with a comprehensive evaluation of diverse methods zhong2020jec. An Italian research group built a new dataset for international private law sovrano2021dataset. A Vietnamese research team (kien-etal-2020-answering) has proposed a legal QA system as well as an annotated corpus for testing their solution.
Romanian Work. For the Romanian language, one particularly useful work is performed in (masala2021jurbert), which proposes jurBERT, a BERT model adapted to the Romanian legal domain through an extensively large corpus of legal text, which amounts to 160GB, according to the authors. They also tested the model on downstream tasks, in which the smaller jurBERT model surprisingly achieved better results. Following this result, the authors attempted to build a judicial prediction system (masala-etal-2024-improving) using the above-mentioned model.
Count Avg Length Max Length
Articles 330,320 95.11 24,735
Words 31,416,577 6.16 28
Vocabulary 78,355 9.59 28
[…]
Edges 319,958 - -
Table 1: General statistics for CROL and KG.
2.2 Graph Based Approaches
The literature features variable approaches when it comes to KG-based QA. We identify 2 main classes of approaches: symbolic and numeric. Since the KG is a symbolic structured representation of factual knowledge, some works revolved around symbolic methods. One work (chakraborty2024multi) explored multi-hop QA approaches using an LLM-KG-based algorithm, which identified the correct answer. The numeric approaches attempt to make use of learned numeric representations of the data to make predictions. One such work is performed in (he-etal-2022-dialmed), which proposes a medical dialogue dataset as well as a method for using a medical graph (wang2019kgat) for predicting the corresponding medication given patient-doctor dialogue. Our work attempts to create a synergy between the two approaches, as well as propose the datasets and the KG for future uses.
3 Novel Resources
Dataset Size Category Language
PrivacyQA (ravichander2019question) 1,750 Span English
PolicyQA (ahmadetal2020policyqa) 714 Span English
JEC-QA (zhong2020jec) 26,367 Multi-Choice Chinese
COLIEE-21 (rabelo2022overview) 887 Binary Japanese
BSARD (louis2021statutory) 1,100 Article Retrieval French
[…]
Table 2: General statistics for CROL and KG.
JuRO. We introduce a new dataset for legal QA, which encompasses 3 different types of examinations: Entrance into the judicial system, entrance into the Bar, and Promotion exams for judicial positions. We are the first to propose such a dataset for the Romanian language and release a Romanian openly available legal dataset. We hope this will open opportunities for the Romanian, multilingual, and low-resource language research. More details for this dataset as well as other resources can be found in Appendix LABEL:appendix:a.
CROL. Corpus for Romanian Law represents a collection of gathered legal documents for all of the law branches mentioned in Figure LABEL:fig:juropie. It can mainly serve as a knowledge base for IR techniques for Romanian legal tasks. There has been past work on creating a Romanian legal corpus, such as the Marcell project (varadi2020marcell), which aimed at creating a multilingual legal corpus, including Romanian law. However, this represents an annotated text corpus useful for NER training in a legal context as well as pertaining, but did not make a distinction between "documents that are in effect and those that are not.” We have made a clear separation between legal documents and their updates.
Law-RoG. We present the first-ever KG for the Romanian language. This KG is built upon the CROL corpus via entity-relation extraction. Particularly, following this work (edge2024local), we use a single prompt for identifying named entities as well as the relations between them and instruct a LLM to output entity-relation-entity triplets in our desired format using in-context (brown2020fewshot) learning abilities that LLMs exhibit via few-shot prompting. We opted for this approach since the Romanian NLP lacks resources for building a specialized pipeline for entity and relation extraction, especially for the legal domain. To make sure this model would generally produce factually and coherently correct information, we sampled 10 documents and their corresponding generated relations from each legal domain, and used a total of 5 human NLP experts, and we all reached a consensus that the outputs were coherent enough and did not hallucinate beyond the given document, and even though not perfect, we concluded they are more than enough for almost all related NLP applications.
3.1 Data Collection
JuRO dataset represented an extensive work for it to reach the final current form. The data was found in multiple forms and had to resort to multiple tools and stages. Where possible, we used web scrapping to extract questions and their corresponding answers. Further, we scrapped PDF documents with textual information available. Lastly, we used OCR to extract the underlying text from the scanned PDFs. In addition, we went for manual verification of data because some scans presented high-level damages of noise, which made OCR impractical, and even though manual extraction was tedious, it was more time economical than going through the OCR process at this stage. All exams are publicly available on the official portals of the institutions responsible for examinations.
The CROL corpus was extracted exclusively via scrapping from the official Ministry of Justice portal, department of the Romanian Government, in a different manner as opposed to the Marcell project (varadi2020marcell), by ensuring an organized crawling of the portal and extracting legal documents along with their distinct legal articles and their updates.
4 GRAF
We introduce a novel approach for retrieving information from a KG, which we applied to legal MCQA. The same principles discussed regarding claim checking and validation can be applied to other tasks requiring factual knowledge.
4.1 Problem Formulation
At first glance, we are presented with questions that may have only 1 correct answer for a dataset, and in other scenarios, a question can have up to 2 correct choices. We will make a distinction between these 2 settings and will formulate the goal of the problem according to the architecture of the proposed model. In any case, the dataset can be represented as $\mathcal{D}=\{e_{i}=(Q_{i},C_{i},T_{i})|i\in\{1..|\mathcal{D}|\}\}$. Each triplet entry $e_{i}$ contains the question body $Q_{i}$ in textual form, the set of possible answer choices $C_{i}=\{C^{1}_{i},C^{2}_{i},C^{3}_{i}\}$ and the set of target answers $T_{i}$ with $|T_{i}|\leq 2$ and $T_{i}\subset C_{i}$. Each answer choice is a tuple $C^{j}_{i}=(\sigma^{j}_{i},\epsilon^{j}_{i})$ where $\sigma^{j}_{i}$ is the symbol associated with the text content $\epsilon^{j}_{i}$ of the answer choice. We will now look at 2 classes of models we also experimented with and formulate the learning goal.
Autoregressive Models. These models are known to exhibit impressive capabilities in generative tasks. They can also be adapted to classification tasks by teaching them the correlation between the class concept and the chosen class symbol. Their goal is to minimize the negative log-likelihood of the class symbol conditioned on the given input. Specifically:
$\mathcal{L}=-\sum\log P(t^{j}_{i}|\mathcal{Z}(Q_{i},C_{i},t^{k}_{i});\boldsymbol{\theta})$ (1)
where $j>k$ and $t^{0}_{i}$ represents the empty sequence. The $\mathcal{Z}$ function maps a given triplet to a sequence that can be processed by the given probabilistic model, known as the model prompt. The prompt has the purpose of facilitating and guiding the model towards a lower on-average negative log-likelihood and consequently, the correct answer. Even though this work also explores sequence-to-sequence models, the ones we chose in particular also generate the output in an autoregressive fashion via the decoder model, and thus, our previous discussion still holds.
Encoder Models. These models showed excellent performances on classification tasks despite their relatively smaller size in practice. They feature a good semantic understanding of a given sequence via their pretraining objectives. For instance, BERT (devlin2018bert) featured word level and sentence level pretraining, which allowed it to gain a semantic understanding of language at these levels. However, they do not exhibit symbol level correlation (robinson2023leveraging) as LLMs do, and thus, we resort to using their semantic understanding of textual sequences for outputting a number that represents the degree to which a given choice is correct given a question. We consider two learning goals for these models, the binary cross entropy minimization for models outputting probabilities:
$\mathcal{L}_{1}=-(o^{j}_{i}\log(y^{j}_{i})+(1-o^{j}_{i})\log(1-y^{j}_{i}))$ (2)
where $o^{j}_{i}=\mathbbm{1}_{T_{i}}(C^{j}_{i})$ and $y^{j}_{i}$ is the model output probability. We also use the cosine similarity embedding loss to align a given question with the correct answer choice:
$\mathcal{L}_{2}=(1+o^{j}_{i})(1-y^{j}_{i})+(1-o^{j}_{i})y^{j}_{i}$ (3)
where $o^{j}_{i}$ has the same meaning as above except the negative class becomes -1 instead of 0, $y^{j}_{i}=\cos(\bar{Q_{i}},\bar{C^{j}_{i}})$ and we refer to the bar notation as the embeddings of the question and choice respectively.
For the inference, we consider the question along with the set of choices and select the top $|T_{i}|$ scores in the following way:
Figure 1: GRAF procedure overview.
$Y_{i}^{*}=\text{TopK}(Y_{i},|T_{i}|)$ (4)
where $Y_{i}=\{y^{j}_{i}|y^{j}_{i}=\text{Score}(Q_{i},C^{j}_{i})\}$. TopK is a generalized argmax function that selects the best K candidates from a given list. In the end, the chosen options by the model are $C^{k}_{i}$ with $k\in Y_{i}^{*}$.
Figure 2: Translated sample question from the JuRO dataset and a showcase of GRAF on this sample. In this case, the entities "Self-defense" and "Unauthorized Access" have the best alignment with the entities in choice "C" thus it is most likely to be the correct answer.
Data: $Q$ Question, $C$ Choices
$G$ Knowledge Graph
Result: $P$ Choice Proababilities
$P\leftarrow\text{[]}$]
$C\leftarrow$ split_choices($C$)
for $c_{i}\in C$ do
        /* Query Claim Extraction model and obtain $CG$ Claim Graph */
       $CG\leftarrow\text{claim\_extract}(Q||c_{i})$
        /* Sample $G$ into Sub Graph $SG$/
       $SG\leftarrow\text{sample\_graph}(Q||c_{i})$
        /* Encode $SG$ and $CG$/
       ${}{}SGE,CGE\leftarrow\text{\{\}},\text{\{\}}$} , {}
       for $(E_{a},r_{ab},E_{b})\in SG$ do
             $SGE\text{.append}(\text{enc}(E_{a}),\text{enc}(r_{ab}),\text{enc}(E_{b}))$
       end for
      for $(E_{a},r_{ab},E_{b})\in CG$ do
             $CGE\text{.append}(\text{enc}(E_{a}),\text{enc}(r_{ab}),\text{enc}(E_{b}))$
       end for
      $SGE,CGE\leftarrow\text{GAT}(SGE),\text{GAT}(CGE)$
       for $(h^{i}_{c},h^{i})\in CGE\times SGE$ do
             $R^{ij}\leftarrow\cos(h^{i}_{c},h_{i})$
       end for
      $\bar{H}\leftarrow Rh$
       $c\leftarrow\text{enc}(Q||C)$
       $c_{\text{final}},H_{\text{final}}\leftarrow\text{SelfAttention}(c||\bar{H})$
       $P\text{.append}(c_{\text{final}})$
end for
Algorithm 1 GRAF
4.2 Algorithm Description
Our algorithm runs for every given question along with the set of choices it presents. In other words, the input to our algorithm is the question-answer pair and the KG.
Claim Graph. A multiple-choice question is mainly made up of choices that present different claims with different truth values. To pick the most suitable answer from a set of choices, we first decompose each (Question, Choice) pair into their underlying claims by using a LLM in our case, but more lightweight solutions can be adopted with the adequate resources. We called this procedure Cross Claim Extraction, as we are interested in claims whose entities come from distinct places from either the question and the given choice, or both entities, could come from the choice as we do not consider the question body itself to present any untrue claims given the format of the exams. Questions generally present hypothetical scenarios or premises that do not contradict the law. In this sense, we instruct the LLM to receive two distinct texts and extract the claims in the aforementioned fashion.
KG Extraction. We first obtain the domain-specific KG from Law-RoG by using the law branch provided in the question. Since it is often unfeasible to consider the entire graph for inference, we resort to sampling the graph via a procedure which will retrieve the most relevant nodes and edges from the KG using a heuristic. In our case, we use a bag-of-words (BoW) approach and incorporate the BM25 retrieval algorithm (robertson1976relevance). Before inputting the data directly to this algorithm, we preprocess the words by tokenizing and lemmatizing them using the model provided in the SpaCy package for the Romanian language to achieve better, less noise results. After selecting the top k entities in the KG, we proceed to select their vicinity with a BFS at a capped depth. Moreover, we also limit the maximum number of entities that are retrieved during this stage. In our work, we use a depth of 1, select the top 10 entities, nodes, and edges from the KG, and select no more than 50 distinct entities.
KG Encoding. To encode the KG, we first use a language encoder model to embed the entities in the KG. After embedding, we use an adapted version of Graph Attention (GAT) (velickovic2018graph) for further capturing the topological information for each entity. Since the original GAT model was developed only for graphs with no relation encoding, we adapted this model for our purpose to work on Knowledge Graphs (KGs). First, we capture relational topological information by computing the attention coefficients for relations in which the current entity $\boldsymbol{e^{ij}_{E}}$ is involved between the current node and the adjacent edges:
$\boldsymbol{e^{ij}_{E}}=\sigma_{A}(\boldsymbol{(a^{E})^{T}}[\boldsymbol{W_{N}h^{i}_{N}}||\boldsymbol{W_{E}h^{j}_{E}}])$ (5)
and calculate the attention coefficients for the nodes to capture inter-entity relations between the current node and neighboring nodes:
$\boldsymbol{e^{ij}_{N}}=\sigma_{A}(\boldsymbol{(a^{N})^{T}}[\boldsymbol{W_{N}h^{i}_{N}}||\boldsymbol{W_{N}h^{j}_{N}}])$ (6)
[…]
We obtain nodes and edges representations by aggregating the information from the adjacent nodes for each node into $\boldsymbol{h^{{}^{\prime}}_{N}}$ and the information from the adjacent edges for each node into $\boldsymbol{h^{{}^{\prime}}_{E}}$:
$\boldsymbol{h^{{}^{\prime}}_{N}}=\text{softmax}(\boldsymbol{e_{N}})\boldsymbol{W_{N}h_{N}}$ (7)
$\boldsymbol{h^{{}^{\prime}}_{E}}=\text{softmax}(\boldsymbol{e_{E}})\boldsymbol{W_{E}h_{E}}$ (8)
In the end, we combine this information into a single representation for each node into $h^{{}^{\prime}}$ as follows:
$\boldsymbol{h^{{}^{\prime}}}=\boldsymbol{h^{{}^{\prime}}_{N}}+\boldsymbol{h^{{}^{\prime}}_{E}}$ (9)
Final Score. After encoding the graphs, we will select the relevant information from the provided knowledge given the encoded claims. For this, we compute a relevance matrix by calculating the alignment between each encoded claim from the previous discussion, which we denote $\boldsymbol{h^{i}_{c}}$ and all the encoded law sub-KG relations $\boldsymbol{h^{j}}$. We calculate the alignment using the cosine similarity in the following way:
$\boldsymbol{R^{ij}}=\cos(\boldsymbol{h^{i}_{c}},\boldsymbol{h^{j}})$ (10)
We then use this matrix to finally aggregate all the relevant information from the law sub-KG into a matrix containing as many vectors as nodes in the original claim graph, each vector being a numeric encoding representation for every encoded claim node, which in turn represents an embedding for all the neighboring claims:
$\boldsymbol{\bar{H}}=\boldsymbol{Rh}$ (11)
We separately encode the (Question, Choice) pair and decide what information is better suited in the final decision for the current choice. We employ a self-attention mechanism for this task and provide a score based on the gathered information and the given choice:
$[\boldsymbol{c_{\text{final}}}||\boldsymbol{H_{\text{final}}}]=\text{SelfAttention}([\boldsymbol{\bar{c}}||\boldsymbol{\bar{H}}])$ (12)
The term $\bar{c}$ is the encoding of the current (Question, Choice) pair. Finally, we compute the score:
$\text{score}=\sigma(\boldsymbol{W_{\text{final}}}\boldsymbol{c_{\text{final}}})$ (13)
where $\sigma$ is the sigmoid logistic activation used to provide a probability score.
5 Baselines
In this section, we will describe each baseline model used in our experiments. For encoder models, we opt for similar approaches to ones in IR and RAG (wang2024utilizing). In Appendix LABEL:sec:appendix, more experimental details are discussed. As baselines we proposed QBERT, ColQBERT (manotumruksa2020crossbert), ColBERT (khattab2020colbert), Large Language Models (LLMs), LLM with RAG (lewis2020retrieval), LLM fine-tuned with Low Rank Adaptation Method (LoRA) (hu2021lora). More details regarding these baselines can be consulted in the Appendix LABEL:appendix:baselines.
Model Civil Penal Proc Civil Proc Penal Administrative Work Family International Comercial Average
QBERT 35.48 38.29 35.82 40.10 36.63 40.40 38.24 39.29 39.39 38.18
CrossQBERT 41.94 36.04 37.31 41.15 41.58 35.35 34.31 42.86 36.36 38.54
ColBERT 44.09 38.29 47.76 37.50 41.58 42.42 41.18 48.81 40.40 42.45
LLM 48.94 40.18 41.00 42.27 46.53 48.48 49.02 52.38 39.39 45.35
LLM RAG 51.06 43.75 38.81 42.78 57.43 56.57 63.73 52.38 57.58 51.57
LLM LFT 45.74 51.79 74.63 48.97 52.48 49.49 54.90 70.24 53.53 55.75
[…]
Table 3: Results for Promotion exams
6 Results
In this section, we will present the results of our extensive experimentation and discuss the obtained results.
6.1 Quantitative Assessment
We compare the baselines with our algorithm, where it makes sense to compare the models. In the case of Promotion exams, with questions with a single correct answer, we can compare all the proposed models as performed in Table 5. Encoder-based solutions do not know apriori how many answers to predict, and thus, we resort to comparing Encoder models and LLM models on exams with more than one possible correct answer separately and provide the results as discussed in Section 4.1. The reported scores are the ones the model would receive on the test, no additional points given for wrong answers. There are slight inconsistencies in the results per law branch because the training was performed on the entire dataset at once. We reported these scores for having a more fine-grained view of how strong each model is in every legal area. Nevertheless, our model still surpasses all of the baselines on most of the legal branches, particularly 6 out of 9. Tables 2.1 and LABEL:tab:encbar show the results for the encoder-based models on the Entrance and Bar examinations respectively. Likewise, Table LABEL:tab:llmentr shows the results for the Entrance examinations and Table LABEL:tab:llmbar for the Bar examinations.
CrossQBERT. As proposed in (manotumruksa2020crossbert), we proceed by taking the question and the entire set of possible choices and concatenate all of them together in the same fashion as above and consider the first 3 separator tokens along with a single FFN with a sigmoid activation function which will output three scores for the same question corresponding to each of the answer choices. In this way, we believe in providing BERT with more context for it to be able to have additional information about the neighboring choices to make a better, more informed decision.
ColBERT. Initially an architecture used for IR (khattab2020colbert), we use it for our problem for its underlying philosophy: aligning textual representations. We thus use a model to encode the question and a model to encode the individual choice, and we use the resulting embeddings to perform cosine similarity.
LLM. We use the LLM’s generalization capabilities (zhao2023survey), having decent performances on tasks in no-data settings and no further training. We perform prompt engineering (chen2023unleashing) and extensively experiment with multiple prompts and, in the end, provide the results for the prompts that provided the best performances.
LLM RAG. We make use of Retrieval Augmented Generation (RAG) gao2023retrieval, to provide LLMs with contextual information that would answer the question or guide the model towards the answer. We use the BM25 retriever (robertson1976relevance) along with the SpaCy package for text normalization (tokenization and lemmatization) to extract the top 10 most relevant documents from the corpus. We take the articles from the corpus and chunk them into 50-word documents. We allow consecutive chunks to overlap by 25 words to maintain the context and not potentially disrupt abruptly the flow of the information presented.
LLM LFT. Finally, for our experiments, we fine-tune these LLMs using the LoRA adaptation method (hu2021lora), which was shown experimentally to match the performances of classic full-parameter fine-tuning. This, together with the previous baseline achieve the best results among the baselines. We opt for LoRA method since our computational resources would not allow a full-parameter fine-tuning of all our proposed LLMs.
Appendix D Models
The BERT models we used in our work are the RoBERT (masala-etal-2020-robert), a Romanian BERT model trained on general domain and jurBERT (masala2021jurbert) a Romanian BERT model trained on legal domain. After attempting multiple LLM models and comparing their preliminary performances on this task, we proposed the best performing LLMs, of reasonable sizes for our resources for the conducted experiments. We used Flan-T5 (raffel2020exploring), XL 3B and XXL 11B variants, and Mistral 7B (jiang2023mistral). These LLMs are instruction tuned, and we opted for this kind of LLMs for their better performances on instruction following and target tasks (wei2022finetuned), and we could leverage their initial performance on further fine-tuning. For the GAT model, we used 6 attention heads as we noticed experimentally that this value represents an equilibrium in our view between the number of average non-zero entries and computational demands when the GAT is initialized and tested given inputs aiming to potentially ameliorate the dying gradient phenomenon from the null entries. For KG construction and claim extraction we used the Mixtral-8x7B-Instruct LLM (jiang2024mixtral) quantized to 8 bits using the int8 algorithm (dettmers20228bit) implemented in the bitsandbytes library. Even though this model is fairly large, we used it in the KG construction and claim extraction process just as a trustworthy mean which will be more likely to correctly extract the entities and the relations. More lightweight solutions can of-course be built by training a smaller language model for this task or by in fact distilling (hinton2015distilling; gu2023knowledge) Mixtral to a Small Language Model (SLM) for instance. Mixtral did not contribute in any way to helping the rest of our framework in taking the right choice, it had a very clear and definite role in our algorithm which can be easily replaced with any other lightweight solution which we could not perform due to the lack of available data in this sense for the Romanian language.
Appendix E Evaluation Metric
As metric, we used the score that a model would receive on the given test, which is equivalent to the accuracy of the model on the task. No extra points are given if the model miss-predicts correct answers or does not include all correct answers. We used this metric to emphasize the actual test performances of the models. Moreover, we argued that the dataset is balanced and actually cared for comparative analysis and thus we considered this metric will suffice in order to not overload the results sections with more numbers.
Appendix F Detailed Evaluation
Tables 14, 14, 14 show all of our extensive evaluation conducted on different backbone models and different prompts. P1 and P2 refer to the best prompts for Mistral and FLAN-T5 models respectively. Our approach surpasses all of the baseline combinations by using jurBERT-large (masala2021jurbert) as our backbone encoder model. Moreover, it is completely above encoder models in all settings.
Figure 6: The token distribution for the JuRo dataset for Flan-T5 and Mistral models.
Task Train/Test/Validation Instances # Classes
Entrance Civil Law 175/50/26 6
Bar Civil Law 432/123/63 6
Promotion Civil Law 326/93/47 3
[…]
Entrance Penal Law 175/50/26 6
Bar Penal Law 430/123/63 6
Promotion Penal Law 775/220/111 3
[…]
Entrance Procedural Civil Law 174/50/26 6
Bar Procedural Civil Law 431/123/63 6
Promotion Procedural Civil Law 207/67/29 3
[…]
Entrance Procedural Penal Law 174/50/26 6
Bar Procedural Penal Law 432/123/63 6
Promotion Procedural Penal Law 676/123/97 3
[…]
Promotion Administrative Law 351/99/51 3
Promotion Commercial Law 344/98/50 3
Promotion Family Law 355/96/52 3
Promotion International Law 289/83/42 3
Promotion Work Law 343/95/50 3
Table 11: A detailed view of the JuRo dataset.
Model Civil Penal Proc Civil Proc Penal Administrative Work Family International Comercial Average
QBERT
RoBERT-small 29.03 38.29 34.33 29.17 30.69 31.31 31.37 27.38 39.39 32.33
RoBERT-base 25.81 37.39 34.33 32.29 36.63 37.37 37.25 27.38 36.36 33.87
RoBERT-large 33.33 35.14 29.85 35.94 27.72 31.31 38.24 30.95 36.36 32.20
jurBERT-base 35.48 28.38 35.82 37.50 25.74 40.40 29.41 28.57 30.30 32.40
jurBERT-large 32.26 34.68 31.34 40.10 25.74 27.27 29.41 39.29 39.39 29.94
CrossQBERT
RoBERT-small 35.48 33.78 32.84 38.54 41.58 26.26 28.43 42.86 35.35 35.01
RoBERT-base 41.94 28.83 28.36 28.65 36.63 33.33 28.43 26.19 30.30 31.41
RoBERT-large 41.94 29.28 34.33 29.69 31.68 29.29 24.51 35.71 35.35 32.42
jurBERT-base 41.94 31.08 37.31 35.42 30.69 35.35 29.41 30.95 36.36 34.28
jurBERT-large 29.03 36.04 35.82 41.15 30.69 35.35 34.31 26.19 34.34 33.66
ColBERT
RoBERT-small 33.33 31.98 32.83 36.46 32.67 33.33 37.25 44.05 34.34 35.14
RoBERT-base 40.86 29.28 31.34 33.33 41.58 31.31 41.18 48.81 35.35 37.00
RoBERT-large 33.33 38.29 37.31 37.50 27.72 29.29 36.27 32.14 40.40 36.69
jurBERT-base 36.56 33.78 37.31 35.42 22.77 42.42 32.35 35.71 33.33 34.41
jurBERT-large 44.09 32.43 47.76 32.81 35.64 36.36 36.27 36.90 39.39 37.96
LLM P1
FLAN-T5 XL 32.98 38.39 39.83 36.08 40.59 34.34 40.20 35.71 39.39 37.50
FLAN-T5 XXL 41.49 35.26 34.12 31.44 34.65 33.33 40.19 29.76 33.33 34.84
Mistral-Instruct v0.1 47.87 35.27 34.96 37.11 44.55 43.43 46.07 41.67 37.37 40.92
Mistral-Instruct v0.2 41.49 39.73 35.63 42.27 43.56 48.48 49.02 47.62 37.37 42.80
LLM P2
FLAN-T5 XL 39.36 39.29 41.00 35.57 39.60 39.39 40.20 45.24 35.35 39.44
FLAN-T5 XXL 48.94 40.18 38.00 37.63 46.53 37.37 48.04 52.38 37.37 42.94
Mistral-Instruct v0.1 41.49 32.14 31.76 33.00 34.65 35.35 28.43 41.67 32.32 34.53
Mistral-Instruct v0.2 38.30 31.70 27.56 34.00 44.55 37.37 37.25 30.95 34.34 35.11
LLM RAG
FLAN-T5 XL 51.06 39.29 32.84 40.21 47.52 47.47 58.82 46.43 51.52 46.13
FLAN-T5 XXL 47.87 43.75 31.34 42.27 55.45 56.57 60.78 47.62 57.58 49.25
Mistral-Instruct v0.1 42.55 34.82 34.33 35.05 51.49 41.41 55.88 47.62 50.51 43.74
Mistral-Instruct v0.2 51.06 39.29 38.81 42.78 57.43 53.54 63.73 45.24 56.57 49.83
LLM LFT
FLAN-T5 XL 45.74 50.00 71.64 43.30 48.51 45.45 54.90 55.95 45.45 51.22
FLAN-T5 XXL 45.74 51.79 74.63 48.97 50.50 49.49 52.94 70.24 53.53 55.31
Mistral-Instruct v0.1 38.30 46.43 71.64 42.78 52.48 49.49 54.90 65.48 47.47 52.11
Mistral-Instruct v0.2 42.55 51.79 67.16 45.88 48.51 47.47 53.92 70.24 49.49 53.00
GRAF(Ours)
RoBERT-small 35.48 49.10 67.69 42.11 50.00 48.96 48.48 54.79 45.26 49.10
RoBERT-base 46.24 50.45 69.23 42.63 45.91 57.29 57.58 58.90 41.05 52.14
RoBERT-large 47.31 52.70 66.15 51.05 59.18 51.04 56.57 61.64 50.53 55.13
jurBERT-base 49.46 49.55 78.46 46.84 53.06 50.00 57.58 58.90 48.42 54.70
jurBERT-large 45.16 47.30 69.23 45.79 47.96 54.17 68.69 67.12 56.84 55.81
Table 12: Results for Promotion exams
Model Civil Penal Proc Civil Proc Penal Average
QBERT
RoBERT-small 29.41 49.10 37.25 33.33 37.27
RoBERT-base 27.45 45.09 27.45 25.49 31.37
RoBERT-large 29.41 37.25 37.25 35.29 34.80
jurBERT-base 41.18 33.33 33.33 31.37 34.80
jurBERT-large 31.37 21.57 33.33 39.22 31.37
CrossQBERT
RoBERT-small 21.57 49.02 43.14 37.25 37.75
RoBERT-base 31.37 27.45 27.45 25.49 27.94
RoBERT-large 41.18 29.41 29.41 31.37 32.84
jurBERT-base 25.49 35.29 35.29 42.14 34.55
jurBERT-large 35.29 41.18 43.14 27.45 36.77
ColBERT
RoBERT-small 23.53 33.33 35.29 27.45 29.90
RoBERT-base 33.33 29.41 27.45 21.57 27.94
RoBERT-large 21.57 43.14 50.98 31.37 36.77
jurBERT-base 31.37 33.33 43.14 41.18 37.26
jurBERT-large 45.10 41.18 29.41 31.37 36.77
LLM P1
FLAN-T5 XL 25.49 21.57 11.76 27.45 21.56
FLAN-T5 XXL 35.29 31.37 13.72 23.53 25.97
Mistral-Instruct v0.1 29.41 13.72 11.76 25.49 20.10
Mistral-Instruct v0.2 26.31 21.57 11.76 21.56 20.29
LLM P2
FLAN-T5 XL 29.41 21.57 13.72 27.45 23.04
FLAN-T5 XXL 31.37 25.49 15.69 25.49 24.51
Mistral-Instruct v0.1 33.33 21.57 9.80 25.49 22.55
Mistral-Instruct v0.2 39.21 19.61 11.76 21.57 23.04
LLM RAG
FLAN-T5 XL 31.37 17.65 21.57 23.53 23.53
FLAN-T5 XXL 39.22 27.45 21.57 27.45 28.92
Mistral-Instruct v0.1 11.76 7.84 5.88 15.69 10.29
Mistral-Instruct v0.2 43.14 23.53 21.57 27.45 28.92
LLM LFT
FLAN-T5 XL 35.29 45.10 41.18 41.18 40.69
FLAN-T5 XXL 37.25 43.14 35.29 29.41 36.27
Mistral-Instruct v0.1 19.61 29.41 29.41 27.45 26.47
Mistral-Instruct v0.2 17.65 17.65 17.65 23.53 19.12
GRAF(Ours)
RoBERT-small 43.14 52.94 45.10 45.10 46.57
RoBERT-base 60.78 58.82 41.18 50.98 52.94
RoBERT-large 52.94 62.75 54.90 56.86 56.86
jurBERT-base 47.06 52.94 50.98 45.10 49.02
jurBERT-large 54.90 54.90 52.94 49.01 52.94
Table 13: Results for Entrance exams
Model Civil Penal Proc Civil Proc Penal Average
QBERT
RoBERT-small 37.90 32.26 33.06 32.26 33.87
RoBERT-base 26.61 27.42 29.84 36.29 30.04
RoBERT-large 34.68 32.26 40.32 33.87 35.28
jurBERT-base 33.87 23.58 33.06 33.06 30.89
jurBERT-large 40.32 31.71 32.26 30.65 33.74
CrossQBERT
RoBERT-small 34.68 31.97 39.52 34.68 35.21
RoBERT-base 30.65 38.52 29.84 30.65 32.42
RoBERT-large 41.94 36.07 33.87 34.68 36.64
jurBERT-base 22.58 34.43 29.84 34.68 30.28
jurBERT-large 36.29 27.05 31.45 30.65 31.36
ColBERT
RoBERT-small 29.03 33.06 28.23 27.42 29.44
RoBERT-base 38.71 27.42 34.68 29.84 32.66
RoBERT-large 36.29 32.26 36.29 39.52 36.09
jurBERT-base 23.39 29.03 31.45 35.48 29.84
jurBERT-large 29.03 33.87 35.48 33.87 33.06
LLM P1
FLAN-T5 XL 19.35 25.00 17.74 12.90 18.75
FLAN-T5 XXL 21.77 24.19 23.38 18.54 21.97
Mistral-Instruct v0.1 14.52 12.10 16.13 13.70 14.11
Mistral-Instruct v0.2 19.35 22.58 15.32 14.52 17.94
LLM P2
FLAN-T5 XL 17.74 22.58 19.35 13.71 18.35
FLAN-T5 XXL 18.55 24.19 23.39 13.71 19.96
Mistral-Instruct v0.1 14.52 18.55 16.13 12.90 15.53
Mistral-Instruct v0.2 15.32 24.19 16.13 14.52 17.54
LLM RAG
FLAN-T5 XL 21.77 22.58 20.16 16.94 20.36
FLAN-T5 XXL 20.97 26.61 33.87 20.97 25.61
Mistral-Instruct v0.1 5.65 11.29 12.90 5.65 8.87
Mistral-Instruct v0.2 26.61 29.03 20.16 19.35 23.79
LLM LFT
FLAN-T5 XL 31.45 32.26 29.03 36.29 32.26
FLAN-T5 XXL 27.42 28.23 33.06 22.58 27.82
Mistral-Instruct v0.1 27.42 31.45 21.77 29.03 27.42
Mistral-Instruct v0.2 13.71 14.52 24.19 23.39 18.95
GRAF(Ours)
RoBERT-small 46.77 39.52 43.90 43.44 43.41
RoBERT-base 48.39 49.19 52.85 46.72 49.29
RoBERT-large 50.00 47.58 49.59 55.73 50.73
jurBERT-base 51.61 55.65 51.22 59.02 54.38
jurBERT-large 51.61 61.29 56.10 57.38 56.60
Table 14: Results for Bar exams.
Appendix G Romanian Prompts
P1 - Entrance, Bar Răspunde la următoarea întrebare de legalitate din {tip drept}. Cel mult 2 răspunsuri sunt corecte. Dacă un singur răspuns este corect, vei răspunde doar cu litera răspunsului corect. Dacă 2 răspunsuri sunt corecte, vei răspunde doar cu literele răspunsurilor corecte: {tip drept} {întrebare} {variante de răspuns}
P1 - Promotion Răspunde la următoarea întrebare de legalitate din {tip drept}. Un singur răspuns este corect. Tu vei răspunde doar cu litera răspunsului corect: {tip drept} {întrebare} {variante de răspuns}
P2 - Entrance, Bar Răspunde la următoarea întrebare de legalitate din {tip drept}. Cel mult 2 răspunsuri sunt corecte. Dacă un singur răspuns este corect, vei răspunde doar cu litera răspunsului corect. Dacă 2 răspunsuri sunt corecte, vei răspunde doar cu literele răspunsurilor corecte Răspunde cu doar unul dintre simbolurile din lista [A, B, C, AB, AC, BC]: {tip drept} {întrebare} {variante de răspuns}
P2 - Promotion Răspunde la următoarea întrebare de legalitate din {tip drept} cu doar una dintre literele din lista [A, B, C]. Un singur răspuns este corect: {tip drept} {întrebare} {variante de răspuns}
[…]
Appendix H Translated Prompts
P1 - Entrance, Bar Answer the following legal question from {law type}. At most 2 answers are correct. If a single answer is correct, you will only answer with the letter of the correct answer. If 2 answers are correct, you will answer with only the letters of the correct answers: {law type} {question} {answer choices}
P1 - Promotion Answer the following legal question from {law type}. A single answer is correct. You will only answer with the letter of the correct answer: {law type} {question} {answer choices}
P2 - Entrance, Bar Answer the following question from {law type}. At most 2 answers are correct. If a single answer is correct, you will only a answer with the letter of the correct answer. If 2 answers are correct, you will answer with the letters of the correct letters. Answer with only one of the symbols from the list [A, B, C, AB, AC, BC]: {law type} {question} {answer choices}
P2 - Promotion Answer the following legal question from {law type} with only one of the letters from the list [A, B, C]. A single answer is correct: {law type} {question} {answer choices}
FLAN-T5 RAG - Entrance, Bar Answer the following legal question from {law type}. At most 2 answers are correct. If a single answer is correct, you will only answer with the letter of the correct answer. If 2 answers are correct, you will answer with only the letters of the correct answers If the information from the context is not in the question, then ignore the context and answer the question. Answer with only one of the symbols from the list [A, B, C, AB, AC, BC]. Context: {documets} Question: {question} {answer choices}
FLAN-T5 RAG - Promotion Answer the following legal question from {law type} with only one of the letters from the list [A, B, C]. If the information from the context is not in the question, then ignore the context and answer the question. A single answer is correct. Context: {documents} Question: {question} {answer choices}
Mistral RAG - Entrance, Bar Answer the following legal question from {law type}. At most 2 answers are correct. If a single answer is correct, you will only answer with the letter of the correct answer. If 2 answers are correct, you will answer with only the letters of the correct answers. If the information from the context is not in the question, then ignore the context and answer the question. Context: {documents} Question: {question} {answer choices}
Mistral RAG - Promotion Answer the following legal question from {law type}. You will only answer with the letter of the correct answer. If the information from the context is not in the question, then ignore the context and answer the question. A single answer is correct. Context: {documents} Question: {question} {answer choices}