### v2

# GRAF: Graph Retrieval Augmented by Facts for Romanian Legal Multi-Choice Question Answering

Abstract
Pre-trained language models have shown remarkable performances in recent years, setting a new paradigm for natural language processing (NLP) research. The legal domain has received some attention from the NLP community partly due to its textual nature. Question-answering (QA) systems represent some of the tasks in this domain. This work explores the legal Multiple-Choice QA (MCQA) for a low-resource language, Romanian. The contribution of this work is multi-fold. We introduce JuRO, the first openly available Romanian legal MCQA dataset, comprising 10,836 questions from three examinations. Along with this dataset, we introduce CROL, an organized corpus of laws with a total of 93 distinct documents with their modifications over 763 time spans, which we used for information retrieval techniques in this work. Additionally, we construct Law-RoG, the first graph of legal knowledge for the Romanian language derived from the corpus mentioned above. Lastly, we propose a novel approach for MCQA, namely Graph Retrieval Augmented by Facts (GRAF), which achieves competitive results with generally accepted state-of-the-art methods and even exceeds them in most settings.
GRAF: Graph Retrieval Augmented by Facts for Romanian Legal Multi-Choice Question Answering
Cristian-George Craciun1,2thanks: cristian.craciun@tum.de Răzvan-Alexandru Smădu1, Dumitru-Clementin Cercel1thanks: dumitru.cercel@upb.ro, Mihaela-Claudia Cercel3,4 1National University of Science and Technology POLITEHNICA Bucharest, Faculty of Automatic Control and Computers, Bucharest, Romania 2Technical University of Munich, Munich, Germany 3Paris 1 Panthéon-Sorbonne University, Paris, France 4University of Bucharest, Bucharest, Romania
footnotetext: Under review.
1 Introduction
Question-answering (QA) represents a family of downstream natural language processing (NLP) tasks explored in various settings (Zhong et al., 2020; Rogers et al., 2023; Singhal et al., 2023). Some QA tasks can be formulated as open-ended questions that require an elaborate answer, which may or may not include the rationale behind it (Chen et al., 2017; Labrak et al., 2022). Other studies focus on retrieving the answer from a given context. In this scenario, the answer can either be found explicit in the provided context, and a model has to identify the location at which the answer occurs (Zaib et al., 2021) or infer the correct answer based on understanding and interpretation of the context at hand, a task known as machine reading comprehension (Baradaran et al., 2022).
The tasks mentioned above have been explored for various domains and languages (Louis and Spanakis, 2022; Hoppe et al., 2021; Askari et al., 2022; Sen et al., 2022; Ekram et al., 2022). Our work focuses on MCQA for the Romanian legal domain. Legal QA represents an emerging area of research due to the insights it can provide in tackling various other problems, such as the dialogue system He et al. (2024), as well as, from a practical perspective, by benefiting the average citizen of a nation with the possibility of finding the answer to a legal inquiry together with experts in the field who will achieve increased productivity.
Additionally, available natural language corpora are scarce for the Romanian language. The only MCQA work was performed by Dima et al. (2024), which proposed a Romanian medical MCQA dataset. Although the Romanian legal domain has been tackled in the past (Masala et al., 2021, 2024), no open-source datasets are available. This motivated us to introduce new comprehensive resources for the Romanian language.
In summary, the contributions of this paper are the following:
  • We release JuROhttps://github.com/craciuncg/JuRO, the first open-source legal dataset available for the Romanian MCQA.
  • We release CROLfootnotemark: , a structured corpus of Romanian law that can be utilized to query the law for answers.
  • We release Law-RoGfootnotemark: , the first knowledge graph for the Romanian legal domain.
  • We present a novel algorithm for the MCQA task called GRAFfootnotemark: .
  • We provide a comprehensive evaluation of our proposed MCQA dataset using state-of-the-art methods and our proposed method.
2 Related Work
2.1 Legal-Domain Question Answering
English QA. Initially, traditional legal QA systems were based on information retrieval techniques, rule-based systems, support vector machines Kim et al. (2014, 2017) and shallow convolutional networks (CNNs) Kim et al. (2015b) to answer legal questions from bar exam questions (Kim et al., 2015a). Later, deep neural networks based on CNNs Do et al. (2017) and BiLSTM John et al. (2017) were used with improved performance. The remarkable achievements of the Transformer models (Vaswani et al., 2017) motivated the use in this domain. Shankar et al. (2023) revolved around privacy-related law using various sources to build their benchmark dataset and evaluated several legal BERT (Devlin et al., 2019) models. Similarly, Hendrycks et al. (2021b) evaluated multiple BERT-based architectures on a newly proposed dataset, showing that the volume of training data and the architecture influence the performance. To support these findings, new datasets were proposed Sovrano et al. (2021), including multidisciplinary large datasets Hendrycks et al. (2021a).
Other Languages. Other languages also took the initiative in the early days. For example, the COLIEE 2015 shared task Kim et al. (2015a) proposed the first legal QA task on the Japanese legal documents provided in both Japanese and English. Bach et al. (2017) proposed an analysis of the Vietnamese transportation laws and evaluated a CRF-based system as a prerequisite stage for QA. An encoder-decoder architecture was proposed by Kien et al. (2020), containing word embeddings, convolution, and attention layers. The model was evaluated on a dataset containing 6,000 Vietnamese legal questions, outperforming existing retrieval-based methods. Later, Vuong et al. (2023) proposed an end-to-end retrieval-based system that used a pre-trained BERT model on weakly labeled data. Zhong et al. (2020) addressed the Chinese legal field, featuring an MCQA dataset containing practice exercise questions, and a knowledge database. The comprehensive evaluation of various methods, such as transformers, attention, and distant supervision, showed a significant gap until the human-level performance was achieved.
2.2 Romanian Legal Domain
[…]
2.3 Knowledge Graph-Based Question Answering
The literature features variable approaches when it comes to KG-based QA. We identify two main classes of approaches: symbolic and numeric. Since KG is a symbolic structured representation of factual knowledge, some works revolved around symbolic methods. Chakraborty (2024) explored multi-hop QA approaches using an LLM-KG-based algorithm, which identified the correct answer. The numeric approaches attempt to employ learned numeric representations of the data to make predictions. One such work is performed by He et al. (2022), which proposes a medical dialogue dataset as well as a method for using a medical graph (Wang et al., 2019) for predicting the corresponding medication given patient-doctor dialogue. Our work attempts to create a synergy between the two approaches and proposes new datasets and a KG for future uses.
3 Novel Resources
Dataset # Examples Category Language Public
PrivacyQA (Ravichander et al., 2019) 1,750 Span English
PolicyQA (Ahmad et al., 2020) 714 Span English
JEC-QA (Zhong et al., 2020) 26,367 Multi-Choice Chinese
COLIEE-21 (Rabelo et al., 2022) 887 Binary Japanese
BSARD (Louis and Spanakis, 2022) 1,100 Article Retrieval French
[…]
Table 1: Comparison of JuRO with other existing datasets.
Corpus Size Category Language Public
Marcell (Váradi et al., 2020) 317k documents, 774M tokens Pretraining Multilingual *
RoJur (Masala et al., 2021) 11M entries, 160GB Cases Romanian
Ostendorff et al. (2020) 200k court rulings Rulings German
Collarana et al. (2018) 62 pages, 64 sections, 24k words Span English
Kien et al. (2020) 8.5k documents, 118k articles Article Vietnamese
[…]
    BG, HR, HU, PL, RO, SK, SL
Table 2: Comparison of CROL with other datasets from various works.
3.1 JuRO
We introduce a new dataset for the legal QA, called JuRO. It contains past examination questions from all of the legal branches examined in Romania. It represents the first dataset of its kind, which we release to the public (see also Table 3).
Dataset Construction. The data is extracted from various official examination portals. Some of the subjects were extracted using OCR, and to avoid errors, we manually inspected all of the data, and thus, the data presents minimal damage. Each entry essentially consists of the body in which a theoretical question regarding a legal aspect is asked and three possible answer choices labeled A, B, and C, out of which at most two answers are correct.
Statistics. The dataset contains questions from three types of examination: entrance into the judicial system, entrance into the bar, and promotion exams for judicial positions. We present the distribution among legal domains and possible answer combinations in Figure 5 from Appendix A. Promotion exams have three possible choices with a single correct answer, whereas the others have up to two possible correct answers. The distribution among correct answers is generally balanced, with a small exception for bar exams. However, it is noteworthy that questions with a single correct answer are predominant and balanced. Data analysis for the JuRO dataset is presented in Appendix A.
JuRO vs Existing Work. We introduce a new dataset for the Romanian legal QA, which encompasses three different types of examinations: entrance, bar, and promotion. We are the first to propose a legal MCQA dataset for the Romanian language and release it to the public. We hope this will open opportunities for future research on Romanian, multilingual, and low-resource language settings. In Table 3, we compare our work against other legal datasets. Although less than half the size of the JEC-QA dataset Zhong et al. (2020), it is larger than other existing datasets for legal QA from other languages.
3.2 CROL
Count Avg Length Max Length
Articles 330,320 95.11 24,735
Words 31,416,577 6.16 28
Vocabulary 78,355 9.59 28
[…]
Edges 319,958 - -
Table 3: General statistics for CROL and Law-RoG KG.
Corpus for Romanian Law (CROL) represents a collection of gathered legal documents for all of the law branches such as civil, penal, work, administration, commercial, family, and international.
Dataset Construction. The CROL corpus construction used the official Ministry of Justice department portalhttps://legislatie.just.ro/ for crawling all of the laws from the covered branches in the JuRO dataset. All resources mentioned above have been extracted from official sources of national state institutions. Therefore, the language is formal and presents little to no grammatical errors.
Statistics. CROL represents a collection of organized legal documents, amounting to 93 distinct laws and 768 different versions of these laws. It contains 330k articles totaling about 31.5M words with a vocabulary of 78.3k words. Statistics are also presented in Table 3.2 and in Appendix A.
CROL vs Existing Work. Our corpus can mainly serve as a knowledge base for information retrieval (IR) techniques for Romanian legal tasks. There has been past work on creating a Romanian legal corpus, such as the Marcell project (Váradi et al., 2020), which aimed at creating a multilingual legal corpus, including Romanian law. However, this represents an annotated text corpus useful for NER training in a legal context as well as pertaining, but does not make a distinction between "documents that are in effect and those that are not.” We have made a clear separation between legal documents and their updates. There are also other efforts Collarana et al. (2018); Kien et al. (2020); Masala et al. (2021); however, they are not available to the public (see Table 3).
3.3 Law-RoG
We introduce the first knowledge graph for the Romanian language, called Law-RoG. This KG is built upon the CROL corpus via entity-relation extraction. Particularly, following the work of Edge et al. (2024), we prompt an LLM to identify named entities and the relations between them to output entity-relation-entity triplets in our desired format using in-context (Brown et al., 2020) learning abilities that LLMs exhibit via few-shot prompting (see Appendices H and I). We opted for this approach since the Romanian NLP lacks resources for building specialized pipelines for entity and relation extraction, specifically for the legal domain. To validate that the LLM produced factually and coherently correct information, we asked 5 human NLP experts to evaluate 10 randomly sampled documents and their corresponding generated relations for each legal domain. They all agreed that the outputs were coherent and did not hallucinate beyond the given document. Although not perfect, we concluded the generated relations were appropriate for almost all related NLP applications. The resulting KG spans 160k nodes and 320k edges (see Table 3.2).
4 GRAF
4.1 Problem Formulation
The multi-choice QA problem can be formulated as follows. Consider a dataset $\mathcal{D}=\{x_{i}=(Q_{i},C_{i},T_{i})\},i=1:\left|\mathcal{D}\right|$, where each triplet entry $x_{i}$ contains the question body $Q_{i}$ in textual form, the set of $|C_{i}|$ possible answer choices $C_{i}=\{C^{k}_{i}\},{k=1:|C_{i}|}$ and the set of target answers $T_{i}$ with $|T_{i}|\leq|C_{i}|$ and $T_{i}\subseteq C_{i}$. Each answer choice is a tuple $C^{j}_{i}=(\sigma^{j}_{i},\epsilon^{j}_{i})$ where $\sigma^{j}_{i}$ is the choice label (e.g., A, B, C) of the answer, and $\epsilon^{j}_{i}$ represents its textual content. In our setting, we look at questions with $|C_{i}|=3$ choice answers and $|T_{i}|\in\{1,2\}$ correct answers, i.e., single-choice and multi-choice QA, depending on the dataset. Moreover, we investigate two classes of models designed to address these QA variations and formulate the learning goal in Appendix B.
Figure 1: GRAF procedure overview.
Figure 2: A showcase of how GRAF works on a sample question from the JuRO dataset translated to English. First, we construct the claim graph using an LLM that extracts the entities (nodes) and relations between them (edges). Based on the sub-KG extracted from Law-RoG and the claim graph, we determine that the entities “Self-defense” and “Unauthorized Access” have the best alignment with the entities in choice “C”, thus it is most likely to be the correct answer.
Data: $Q$ question, $C$ choices, $G$ knowledge graph
Result: $P$ choice probabilities
1 $P\leftarrow\text{[]}$]
2 for $c_{i}\in split\_choices(C)$ do
       // Query the cross claim extraction model and obtain the claim graph CG
3       $CG\leftarrow\text{cross\_claim\_extract}(Q||c_{i})$
       // Sample a subgraph $SG$ from $G$
4       $SG\leftarrow\text{sample\_graph}(G)$
       // Encode $SG$’s components
5       ${}{}SGE,CGE\leftarrow\text{\{\}},\text{\{\}}$} , {}
6       for $(E_{a},r_{ab},E_{b})\in SG$ do
7             $SGE\text{.append}(\text{enc}(E_{a}),\text{enc}(r_{ab}),\text{enc}(E_{b}))$
8       end for
      // Encode $CG$’s components
9       for $(E_{a},r_{ab},E_{b})\in CG$ do
10             $CGE\text{.append}(\text{enc}(E_{a}),\text{enc}(r_{ab}),\text{enc}(E_{b}))$
11       end for
      // Compute the alignment between encoded claims and all encoded relations
12       for $(h^{i}_{c},h^{j})\in\text{GAT}(SGE)\times\text{GAT}(CGE)$ do
13             $R^{ij}\leftarrow\cos(h^{i}_{c},h^{j})$
14
15       end for
      // Compute the embedding for all neighboring claims
16       $\bar{H}\leftarrow Rh$
       // Compute the final score using self-attention
17       $c\leftarrow\text{enc}(Q||C)$
18       $c_{\text{final}},H_{\text{final}}\leftarrow\text{SelfAttention}(c||\bar{H})$
       // Save the score of the current choice
19       $P\text{.append}(c_{\text{final}})$
20 end for
Algorithm 1 GRAF
4.2 Algorithm Description
Our proposed algorithm is applied for every given question and each of its choice answers. Concretely, the inputs to GRAF are the question-choice pair and the KG containing entities as nodes and relationships between entities as edges. The algorithm is shown in Figure 1 and Algorithm 1 and will be presented in what follows.
Claim Graph. A multiple-choice question is mainly composed of choices that present different claims with different truth values. We are interested in determining claims whose entities come from (1) either the question and the given choice or (2) both come from the choice. Since, in our setting, questions generally present hypothetical scenarios or premises that do not contradict the law, we do not consider the question body itself to present any untrue claims. Therefore, to pick the most suitable answer, we decompose each (question, choice) pair into their underlying claims using an LLM, similar to Edge et al. (2024); however, any lighter solutions can be adopted with adequate resources. We call this procedure Cross Claim Extraction. An example is presented in Figure 2.
KG Sampling. We obtain the domain-specific KG (in our setting, from Law-RoG) by using the law branch related to the question. Since it is often unfeasible to consider the entire graph for inference due to its size, we resort to sampling the graph via a procedure to retrieve the most relevant nodes and edges using a heuristic. First, we preprocess the words by tokenizing and lemmatizing them using the SpaCy package for the Romanian language to achieve better, less noisy results. Then, we use a bag-of-words (BoW) approach and incorporate the BM25 retriever (Robertson and Jones, 1976) to select the top k entities in the KG. We proceed to select their vicinity with a breadth-first search for a limited depth. We also limit the maximum number of entities retrieved during this stage. In our work, we use a depth of 1, select the top 10 entities, nodes, and edges from the KG, and select no more than 50 distinct entities.
KG Encoding. To encode the KG, we use a language encoder model to embed the entities and the relations resulting in the set of node and edge features $\boldsymbol{h_{N}}$ and $\boldsymbol{h_{E}}$, respectively. Then, we adapt the Graph Attention (GAT) (Veličković et al., 2018) to further capture the topological information for each entity. The original GAT model was developed only for graphs with no relation encoding. Therefore, we transform the set of features considering two different linear transformations parametrized by shared $\boldsymbol{W_{N}}$ and $\boldsymbol{W_{E}}$ for the nodes and edges, respectively. To capture the relational topological information, we compute the attention coefficients for relations $\boldsymbol{e^{ij}_{E}}$ in which the current entity is involved between the current node $i$ and the adjacent edges $j$:
$\boldsymbol{e^{ij}_{E}}=\sigma_{A}(\boldsymbol{(a_{E})^{T}}[\boldsymbol{W_{N}h^{i}_{N}}||\boldsymbol{W_{E}h^{j}_{E}}])$ (1)
and calculate the attention coefficients for the nodes $e^{ij}_{N}$ to capture inter-entity relations between the current node $i$ and neighboring nodes $j$:
$\boldsymbol{e^{ij}_{N}}=\sigma_{A}(\boldsymbol{(a_{N})^{T}}[\boldsymbol{W_{N}h^{i}_{N}}||\boldsymbol{W_{N}h^{j}_{N}}])$ (2)
where $\boldsymbol{a_{N}}$ and $\boldsymbol{a_{E}}$ represent two distinct attention vectors for nodes and edges, respectively. We also use a nonlinearity $\sigma_{A}$ as Veličković et al. (2018), which in our case is the LeakyReLU activation function. The $\cdot^{T}$ operator represents transposition, while ||||| | is the concatenation operator.
We obtain final nodes and edges representations by aggregating the information from the adjacent nodes for each node into $\boldsymbol{h^{{}^{\prime}}_{N}}$ and the information from the adjacent edges for each node into $\boldsymbol{h^{{}^{\prime}}_{E}}$:
$\boldsymbol{h^{{}^{\prime}}_{N}}=\text{softmax}(\boldsymbol{e_{N}})\boldsymbol{W_{N}h_{N}}$ (3)
$\boldsymbol{h^{{}^{\prime}}_{E}}=\text{softmax}(\boldsymbol{e_{E}})\boldsymbol{W_{E}h_{E}}$ (4)
In the end, we combine this information into a single representation for each node into $h^{{}^{\prime}}$ as follows:
$\boldsymbol{h^{{}^{\prime}}}=\boldsymbol{h^{{}^{\prime}}_{N}}+\boldsymbol{h^{{}^{\prime}}_{E}}$ (5)
Final Score. After encoding the graphs, we select the relevant information from the provided knowledge given the encoded claims. For this, we compute a relevance matrix by calculating the alignment between each encoded claim $\boldsymbol{h^{i}_{c}}$ and all the encoded law sub-KG relations $\boldsymbol{h^{j}}$. We calculate the alignment using the cosine similarity:
$\boldsymbol{R^{ij}}=\cos(\boldsymbol{h^{i}_{c}},\boldsymbol{h^{j}})$ (6)
We then use this matrix to finally aggregate all the relevant information from the law sub-KG into a matrix containing as many vectors as nodes in the original claim graph, each vector being a numeric encoding representation for every encoded claim node, which in turn represents an embedding for all the neighboring claims:
$\boldsymbol{\bar{H}}=\boldsymbol{Rh}$ (7)
We separately encode the (question, choice) pair into $\bar{c}$ and decide what information is better suited in the final decision for the current choice. We employ a self-attention mechanism for this task and provide a score based on the gathered information and the given choice:
$[\boldsymbol{c_{\text{final}}}||\boldsymbol{H_{\text{final}}}]=\text{SelfAttention}([\boldsymbol{\bar{c}}||\boldsymbol{\bar{H}}])$ (8)
Finally, we compute the score:
$\text{score}=\sigma(\boldsymbol{W_{\text{final}}}\boldsymbol{c_{\text{final}}})$ (9)
where $\sigma$ is the sigmoid logistic activation used to provide a probability score, and $W_{final}$ is a learnable parameter.
5 Experiments and Results
5.1 Baselines
In this section, we present the baseline models used in our experiments. For encoder models, we opt for similar approaches to ones in information retrieval (IR) and retrieval augmented generation (RAG) (Wang et al., 2024). In Appendix C, more experimental details are discussed. As baselines, we proposed QBERT, ColQBERT (Manotumruksa et al., 2020), ColBERT (Khattab and Zaharia, 2020), Large Language Models (LLMs), LLM with RAG (Lewis et al., 2020), and LLM fine-tuned with Low-Rank Adaptation Method (LoRA) (Hu et al., 2021). More details regarding these baselines can be consulted in Appendix D, and regarding the language models employed, check Appendix E. The results are presented in terms of test accuracy which is further detailed in Appendix F.
Model Civil Penal Proc Civil Proc Penal Administrative Work Family International Commercial Average
QBERT 35.48 38.29 35.82 40.10 36.63 40.40 38.24 39.29 39.39 38.18
CrossQBERT 41.94 36.04 37.31 41.15 41.58 35.35 34.31 42.86 36.36 38.54
ColBERT 44.09 38.29 47.76 37.50 41.58 42.42 41.18 48.81 40.40 42.45
LLM 48.94 40.18 41.00 42.27 46.53 48.48 49.02 52.38 39.39 45.35
LLM RAG 51.06 43.75 38.81 42.78 57.43 56.57 63.73 52.38 57.58 51.57
LLM LFT 45.74 51.79 74.63 48.97 52.48 49.49 54.90 70.24 53.53 55.75
[…]
Table 4: Accuracy results for promotion exams.
5.2 Quantitative Assessment
We compare the baselines with our algorithm, where it makes sense to compare the models. In the case of promotion exams, with questions with a single correct answer, we can compare all the proposed models as performed in Table 5.1. Encoder-based solutions do not know how many answers to predict a priori. Thus, we resort to comparing encoder models and LLM models on exams with more than one possible correct answer separately and provide the results as discussed in §4.1. The reported scores are the ones the model would receive on the test, no additional points given for wrong answers. There are slight inconsistencies in the results per law branch because the training was performed on the entire dataset at once. We reported these scores for having a more fine-grained view of how strong each model is in every legal area. Nevertheless, our model still surpasses the baselines on 6 out of 9 legal branches. Tables 5.2 and 5.2 show the results for the encoder-based models on the entrance and bar examinations, respectively. Likewise, Table 7 shows the results for the entrance examinations and Table 8 for the bar examinations. More granular results are presented in Appendix G.
Model Civil Penal Proc Civil Proc Penal Average
QBERT 41.18 49.10 37.25 39.22 41.69
CrossQBERT 41.18 49.02 43.14 37.25 42.65
ColBERT 45.10 43.14 50.98 41.18 45.10
[…]
Table 5: Accuracy results on encoder models for entrance exams.
Model Civil Penal Proc Civil Proc Penal Average
QBERT 40.32 32.26 40.32 36.29 37.30
CrossQBERT 41.94 38.52 39.52 34.68 38.67
ColBERT 38.71 33.87 36.29 39.52 37.10
[…]
Table 6: Accuracy results on encoder models for bar exams.
Model Civil Penal Proc Civil Proc Penal Average
LLM 39.21 31.37 15.69 27.45 28.43
LLM RAG 43.14 27.45 21.57 27.45 29.90
LLM LFT 37.25 45.10 41.18 41.18 41.18
Table 7: Accuracy results on LLMs for entrance exams.
Model Civil Penal Proc Civil Proc Penal Average
LLM 21.77 25.00 23.39 18.54 22.18
LLM RAG 26.61 29.03 33.87 20.97 27.62
LLM LFT 31.45 32.26 33.06 36.29 33.27
Table 8: Accuracy results on LLMs for bar exams.
Figure 3: Comparative performances of different backbone encoder models (i.e., BERT-based models) of different sizes and pre-trained on different domains.
5.3 Qualitative Assessment
We provide a comparative study for the promotion exams to improve our approach over the baselines shown in Figure 4. In Figure 3, we depict the achieved improvements in our approach when increasing the size of the backbone language encoder model and when changing from the language model trained on general topics to a model focused on our target legal domain. Pre-trained models on the legal domain achieve improved performances over their general domain pre-trained counterparts. As seen in Table 5.1, our model clearly surpasses all the baselines, including larger ones in size, i.e., the LLMs. We notice that the fine-tuned LLMs tend to pick some particular law branches very well but lag behind RAG in other areas. GRAF represents the best of both worlds and achieves good performance where training allows and also maintains competitive performance in tasks where RAG excels in fine-tuning. Our model clearly surpasses the baselines from the same class, the encoder models, in every setting and every exam. Moreover, it benefits from the knowledge provided in the KG, the claim-aware IR, and the network structure of the extracted claims. Our procedure captures key information from the given question and assesses the relevance of the interconnections between the entities present in the question and law as illustrated in Figure 2. A single LLM would have to come up internally with such assessments, and architecturally, it is not necessarily designed in such a fashion. In addition, providing context to LLMs has the potential of awakening the hallucination phenomenon and the model might not correctly identify which claims are indeed true given the related legal texts.
Model Avg. Accuracy
GRAF 55.61
   w/o claim graph 53.14↓2.47
   w/o KG 53.61↓2.00
   w/o claim graph + emb. sum 49.51↓6.10
Table 9: Results for different ablated components of our framework.
Figure 4: Comparative performances of best baselines and our approach on different law branches.
5.4 Ablation Study
In Table 9, we present the effect of removing different parts of our algorithm in an ablation study. We eliminate the claim graph and the KG, and we experiment with collapsing the KG embeddings via summation (denoted emb. sum). In this sense, we experiment with various ablated components and found that missing KG or claim graphs underperform the LLM baselines. Even though the model uses the KG, the performances are poorer because there could be irrelevant information that would play the role of distractors in the decision-making process. We show that our claim-aware information extraction mechanism is indeed effective for improved retrieval capabilities and, consequently, performances. Merely collapsing the extracted information via summation would still not make up for the missing claim-aware IR but would completely damage the model’s performance.
6 Conclusions
We have introduced new resources for the Romanian legal domain, being the first to construct and release a Romanian legal KG. We also proposed a novel approach for KG IR, namely GRAF, which surpassed the baselines proposed for MCQA.
Future Work. We expect solutions that will improve upon our proposed method in every possible aspect. Also, there might be solutions that could potentially explore dataset augmentation via LLMs. Studies could be conducted on target-domain IR, which may include multiple languages and JuRO, CROL and even the KG Law-RoG could represent good foundational resources. We hope that our work will motivate more exploration of under-represented languages with low resources and thus would motivate finding solutions that would work in low-resource settings in any language.
Limitations
We have released a legal MCQA dataset by gathering questions from all available law examinations across the country, which has enough samples for training but might not be sufficient for training on one individual law branch, which is the reason why we opted for training on the whole dataset.
The goal of our work is to improve the resources and provide a methodological approach to addressing legal question-answering. Since such systems are meant to help users understand the law, they are not yet entirely accurate. In fact, the best average score achieved by our approach is only 60%. Therefore, more research is required in this direction since the legal domain is a sensitive topic when considering machine learning systems in assessing the laws. We want to acknowledge that, considering the current state, deploying such systems would require human validation from legal experts to reduce any risk of providing unlawful responses.
Ethical Considerations
We have collected our dataset from diverse official public portals. To protect this dataset from inadequate use, we decided to license the use of this dataset only for research purposes and under no circumstances should it be used in commercial settings. Our work was performed so as not to rely on external human crowd-workers and to not raise any ethical concerns. The data does not contain any sensitive personal information regarding any identifiable real person. All of the hypothetically presented scenarios do not use any person’s name and instead use anonymized abbreviations. Since the data was collected from the public domain and made available by applicable law by the administrative institutions in question, we release these resources under the CC BY-NC-SA 4.0 licensehttps://creativecommons.org/licenses/by-nc-sa/4.0/, allowed by the active European regulationshttps://eur-lex.europa.eu/eli/dir/2019/790/oj.
Appendix A Dataset Analysis
Figure 5: On the left, the number of samples from each law category in the JuRO dataset. On the right, the class equilibrium is depicted via color variations in the heatmap. The heatmap scores are normalized by dividing each value by the maximum of the respective row.
The domain distribution of the JuRO, along with the distribution of the answers, are presented in Figure 5. Because of the examinations’ format, at most two answers are correct. However, in the case of promotion exams, only one answer is correct. The domains of the questions are procedural civil, procedural penal, penal, civil, work, administration, commercial, family, and international.
Table 10 presents the TF-IDF scores Salton et al. (1975) for JuRO dataset, calculated using the following formula:
$\text{score}(t,C)=\frac{f(t,C)}{|\{w|d\in C,w\in d\}|}\log\frac{|C|}{|\{d|d\in C\ and\ t\in d\}|}$
where:
  • the current term for which we compute the score is denoted by $t$;
  • $C$ is the corpus of documents, each document containing multiple words;
  • $f(t,C)$ is the frequency of the term $t$ relative to the corpus $C$.
We notice a high score for the word “penal” over other words, indicating a possible prevalence of penal-related content in the dataset. Moreover, the terms “case”, “term”, “appeal”, “judgement”, “court”, and “request” indicate procedures.
Word Translation TF-IDF Score
cerere request 0.03507
instanţă court 0.03352
caz case 0.02769
lege law 0.02627
penal penal 0.02622
hotărâre decision 0.02614
termen term 0.02602
apel appeal 0.02565
sine self 0.02553
judecată judgement 0.02486
Table 10: TF-IDF scores for the top ten words in the JuRO dataset.
In Table 11, we report the TF-IDF scores for the CROL corpus. Generally, there are words commonly found in articles, but no word indicates a significant bias towards some specific law area.
Word Translation TF-IDF Score
articol article 0.03262
caz case 0.02550
persoană person 0.02510
lege law 0.02471
bun good 0.02429
alineat paragraph 0.02314
prevedea stipulate 0.02260
drept just/correct/law 0.02213
număr number 0.01965
public public 0.01961
Table 11: TF-IDF Scores for the top ten words from the CROL corpus.
In Figure 6, we depict the tokens distribution for both FLAN-T5 (Raffel et al., 2020) and Mistral (Jiang et al., 2023) models using their tokenizers. The distributions behave approximately the same; the difference is that the Mistral tokenizer tends to use more tokens to represent the text than the FLAN-T5 one. Table 12 shows a detailed distribution of questions for each examination.
Appendix B Model Architectures
Autoregressive Models. These models are known to exhibit impressive capabilities in generative tasks. They can also be adapted to classification tasks by teaching them the correlation between the class concept and the chosen class symbol. Their goal is to minimize the negative log-likelihood of the class symbol conditioned on the given input. Specifically:
$\mathcal{L}=-\sum\log P(t^{j}_{i}|\mathcal{Z}(Q_{i},C_{i},t^{k}_{i});\boldsymbol{\theta})$ (11)
where $j>k$ and $t^{0}_{i}$ represents the empty sequence. The $\mathcal{Z}$ function maps a given triplet to a sequence that can be processed by the given probabilistic model, known as the model prompt. The prompt has the purpose of facilitating and guiding the model towards a lower on-average negative log-likelihood and consequently, the correct answer. Even though our work also explores sequence-to-sequence models, the ones we chose in particular generate the output in an autoregressive fashion via the decoder module, and thus, our previous discussion still holds.
Encoder Models. These models showed excellent performances on classification tasks despite their relatively smaller size in practice. They feature a good semantic understanding of a given sequence via their pretraining objectives. For instance, BERT (Devlin et al., 2019) featured word-level and sentence-level pretraining, which allowed it to gain a semantic understanding of language. However, they do not exhibit symbol level correlation (Robinson et al., 2023) as LLMs do, and thus, we resort to using their semantic understanding of textual sequences for outputting a number that represents the degree to which a given choice is correct given a question. We consider two learning goals for these models, the binary cross entropy minimization for models outputting probabilities:
$\mathcal{L}_{1}=-(o^{j}_{i}\log(y^{j}_{i})+(1-o^{j}_{i})\log(1-y^{j}_{i}))$ (12)
where $o^{j}_{i}=\mathbbm{1}_{T_{i}}(C^{j}_{i})$ is the ground truth and $y^{j}_{i}$ is the model output probability. We also use the cosine similarity embedding loss to align a given question with the correct answer choice:
$\mathcal{L}_{2}=(1+o^{j}_{i})(1-y^{j}_{i})+(1-o^{j}_{i})y^{j}_{i}$ (13)
where $o^{j}_{i}$ has the same meaning as above except the negative class becomes -1 instead of 0, $y^{j}_{i}=\cos(\bar{Q_{i}},\bar{C^{j}_{i}})$ and we refer to the bar notation as the embeddings of the question and choice respectively.
During inference, we consider the question along with the set of choices and select the top $|T_{i}|$ scores in the following way:
$Y_{i}^{*}=\text{TopK}(Y_{i},|T_{i}|)$ (14)
where $Y_{i}=\{y^{j}_{i}|y^{j}_{i}=\text{Score}(Q_{i},C^{j}_{i})\}$. TopK is a generalized argmax function that selects the best K candidates from a given list. In the end, the chosen options by the model are $C^{k}_{i}$ with $k\in Y_{i}^{*}$.
Appendix C Experimental Setup
Training was performed on the entire JuRO dataset for every model, and for testing, we considered the checkpoint with the best evaluation results during the training phase. For encoders, we use BERT-based models that were trained for 50 epochs even though in almost all cases, the best model was found around epoch 10; we used a learning rate of $10^{-7}$ and the AdamW Kingma and Ba (2015) optimizer via vanilla PyTorch. All BERT models were fine-tuned on all parameters. LLMs were fine-tuned for 50 epochs using the Trainer API provided by the transformers library using a learning rate of $10^{-7}$, AdamW optimizer, LoRA Hu et al. (2022) alpha of 32, LoRA rank 64, and 2 warm-up steps. All our experiments were performed on a single NVIDIA A100 80GB to which we had limited and restricted access. We report the results on a single run.
Appendix D Baselines
QBERT. We consider the BERT model (Devlin et al., 2019) and construct the input to the LM by appending the given question and the choice in the following way: [CLS] + question + [SEP] + choice + [SEP]. We then use the classification token to attach a fast forward network (FFN) on top with a sigmoid activation function, which will report a score between 0 and 1 for the correctness of the answer choice.
CrossQBERT. As proposed by Manotumruksa et al. (2020), we proceed by taking the question and the entire set of possible choices and concatenating them together in the same fashion as for QBERT. We consider the first 3 separator tokens and a single FFN with a sigmoid activation function, which outputs three scores for the same question corresponding to each answer choice. In this way, we provide BERT with more context to have additional information about the neighboring choices to make a better, more informed decision.
ColBERT. Initially, an architecture used for IR (Khattab and Zaharia, 2020), we use it for our task for its underlying philosophy: aligning textual representations. We thus use a model to encode the question and a model to encode the individual choice, and we use the resulting embeddings to perform cosine similarity.
LLM. We use the LLM’s generalization capabilities (Zhao et al., 2023), having decent performances on tasks in no-data settings and no further training. We perform prompt engineering (Chen et al., 2023) and extensively experiment with multiple prompts and, in the end, provide the results for the prompts that provided the best performances. For prompts, see Appendix H and the translations in Appendix I.
LLM RAG. We make use of Retrieval Augmented Generation (RAG) Gao et al. (2023) to provide LLMs with contextual information that would answer the question or guide the model towards the answer. We use the BM25 retriever (Robertson and Jones, 1976) along with the SpaCy package for text normalization (tokenization and lemmatization) to extract the top 10 most relevant documents from the corpus. We take the articles from the CROL corpus and chunk them into 50-word documents. We allow consecutive chunks to overlap by 25 words to maintain the context and not potentially abruptly disrupt the flow of the information presented.
LLM LFT. Finally, for our experiments, we fine-tune these LLMs using the LoRA adaptation method (Hu et al., 2021), which was shown experimentally to match the performances of classic full-parameter fine-tuning. This, together with the previous baseline, achieves the best results among the baselines. We opt for the LoRA method since our computational resources would not allow a full-parameter fine-tuning of all our proposed LLMs.
Appendix E Language Models
The BERT models we used in our work are the RoBERT (Masala et al., 2020), a Romanian BERT model trained on general domain and jurBERT (Masala et al., 2021) – a Romanian BERT model trained on the legal domain. After attempting multiple LLM models and comparing their preliminary performances on this task, we found the best-performing LLMs, of reasonable sizes for our resources for the conducted experiments. We used Flan-T5 (Raffel et al., 2020), XL 3B and XXL 11B variants, and Mistral 7B (Jiang et al., 2023). These LLMs are instruction tuned, and we opted for this kind of LLMs for their better performances on instruction following and target tasks (Wei et al., 2022). We could leverage their initial performance on further fine-tuning. For the GAT model, we used 6 attention heads as we noticed experimentally that this value represents an equilibrium in our view between the number of average non-zero entries and computational demands when the GAT is initialized and tested given inputs aiming to potentially ameliorate the dying gradient phenomenon from the null entries. For KG construction and claim extraction, we used the Mixtral-8x7B-Instruct LLM (Jiang et al., 2024) quantized to 8 bits using the int8 algorithm (Dettmers et al., 2022) implemented in the bitsandbytes library. Even though this model is relatively large, we used it in the KG construction and claim extraction process just as a trustworthy mean which will be more likely to correctly extract the entities and the relations. More lightweight solutions can be built by training a smaller language model for this task or by distilling (Hinton, 2015; Gu et al., 2023) Mixtral to a small language model (SLM), for instance. Mixtral did not contribute in any way to helping the rest of our framework in making the right choice. It had a very clear and definite role in our algorithm, which can be easily replaced with any other lightweight solution which we could not perform due to the lack of available data in this sense for the Romanian language.
Appendix F Evaluation Metric
As for the evaluation metric, we used the score that a model would receive on the given test, equivalent to the model’s accuracy on the task. No extra points are given or taken if the model miss-predicts correct answers or does not include all correct answers. We used this metric to emphasize the actual test performances of the models. Moreover, we argued that the dataset is balanced and cared for comparative analysis, and thus, we considered this metric to suffice to not overload the results sections with more numbers.
Appendix G Detailed Evaluation
Tables 15, 15, 15 show all of our extensive evaluations conducted on different backbone models and different prompts. P1 and P2 refer to the best prompts for Mistral and FLAN-T5 models, respectively. Our approach surpasses all baseline combinations by using jurBERT-large (Masala et al., 2021) as our backbone encoder model. Moreover, it is completely above encoder models in all settings.
Figure 6: The token distribution for the JuRO dataset for Flan-T5 and Mistral tokenizers.
Task Train/Test/Validation Instances # Classes
Entrance Civil Law 175/50/26 6
Bar Civil Law 432/123/63 6
Promotion Civil Law 326/93/47 3
[…]
Entrance Penal Law 175/50/26 6
Bar Penal Law 430/123/63 6
Promotion Penal Law 775/220/111 3
[…]
Entrance Procedural Civil Law 174/50/26 6
Bar Procedural Civil Law 431/123/63 6
Promotion Procedural Civil Law 207/67/29 3
[…]
Entrance Procedural Penal Law 174/50/26 6
Bar Procedural Penal Law 432/123/63 6
Promotion Procedural Penal Law 676/123/97 3
[…]
Promotion Administrative Law 351/99/51 3
Promotion Commercial Law 344/98/50 3
Promotion Family Law 355/96/52 3
Promotion International Law 289/83/42 3
Promotion Work Law 343/95/50 3
Table 12: A detailed view of the JuRO dataset regarding the distribution among legal domains and split into train/test/validation sets.
Model Civil Penal Proc Civil Proc Penal Administrative Work Family International Comercial Average
QBERT
RoBERT-small 29.03 38.29 34.33 29.17 30.69 31.31 31.37 27.38 39.39 32.33
RoBERT-base 25.81 37.39 34.33 32.29 36.63 37.37 37.25 27.38 36.36 33.87
RoBERT-large 33.33 35.14 29.85 35.94 27.72 31.31 38.24 30.95 36.36 32.20
jurBERT-base 35.48 28.38 35.82 37.50 25.74 40.40 29.41 28.57 30.30 32.40
jurBERT-large 32.26 34.68 31.34 40.10 25.74 27.27 29.41 39.29 39.39 29.94
CrossQBERT
RoBERT-small 35.48 33.78 32.84 38.54 41.58 26.26 28.43 42.86 35.35 35.01
RoBERT-base 41.94 28.83 28.36 28.65 36.63 33.33 28.43 26.19 30.30 31.41
RoBERT-large 41.94 29.28 34.33 29.69 31.68 29.29 24.51 35.71 35.35 32.42
jurBERT-base 41.94 31.08 37.31 35.42 30.69 35.35 29.41 30.95 36.36 34.28
jurBERT-large 29.03 36.04 35.82 41.15 30.69 35.35 34.31 26.19 34.34 33.66
ColBERT
RoBERT-small 33.33 31.98 32.83 36.46 32.67 33.33 37.25 44.05 34.34 35.14
RoBERT-base 40.86 29.28 31.34 33.33 41.58 31.31 41.18 48.81 35.35 37.00
RoBERT-large 33.33 38.29 37.31 37.50 27.72 29.29 36.27 32.14 40.40 36.69
jurBERT-base 36.56 33.78 37.31 35.42 22.77 42.42 32.35 35.71 33.33 34.41
jurBERT-large 44.09 32.43 47.76 32.81 35.64 36.36 36.27 36.90 39.39 37.96
LLM P1
FLAN-T5 XL 32.98 38.39 39.83 36.08 40.59 34.34 40.20 35.71 39.39 37.50
FLAN-T5 XXL 41.49 35.26 34.12 31.44 34.65 33.33 40.19 29.76 33.33 34.84
Mistral-Instruct v0.1 47.87 35.27 34.96 37.11 44.55 43.43 46.07 41.67 37.37 40.92
Mistral-Instruct v0.2 41.49 39.73 35.63 42.27 43.56 48.48 49.02 47.62 37.37 42.80
LLM P2
FLAN-T5 XL 39.36 39.29 41.00 35.57 39.60 39.39 40.20 45.24 35.35 39.44
FLAN-T5 XXL 48.94 40.18 38.00 37.63 46.53 37.37 48.04 52.38 37.37 42.94
Mistral-Instruct v0.1 41.49 32.14 31.76 33.00 34.65 35.35 28.43 41.67 32.32 34.53
Mistral-Instruct v0.2 38.30 31.70 27.56 34.00 44.55 37.37 37.25 30.95 34.34 35.11
LLM RAG
FLAN-T5 XL 51.06 39.29 32.84 40.21 47.52 47.47 58.82 46.43 51.52 46.13
FLAN-T5 XXL 47.87 43.75 31.34 42.27 55.45 56.57 60.78 47.62 57.58 49.25
Mistral-Instruct v0.1 42.55 34.82 34.33 35.05 51.49 41.41 55.88 47.62 50.51 43.74
Mistral-Instruct v0.2 51.06 39.29 38.81 42.78 57.43 53.54 63.73 45.24 56.57 49.83
LLM LFT
FLAN-T5 XL 45.74 50.00 71.64 43.30 48.51 45.45 54.90 55.95 45.45 51.22
FLAN-T5 XXL 45.74 51.79 74.63 48.97 50.50 49.49 52.94 70.24 53.53 55.31
Mistral-Instruct v0.1 38.30 46.43 71.64 42.78 52.48 49.49 54.90 65.48 47.47 52.11
Mistral-Instruct v0.2 42.55 51.79 67.16 45.88 48.51 47.47 53.92 70.24 49.49 53.00
GRAF(Ours)
RoBERT-small 35.48 49.10 67.69 42.11 50.00 48.96 48.48 54.79 45.26 49.10
RoBERT-base 46.24 50.45 69.23 42.63 45.91 57.29 57.58 58.90 41.05 52.14
RoBERT-large 47.31 52.70 66.15 51.05 59.18 51.04 56.57 61.64 50.53 55.13
jurBERT-base 49.46 49.55 78.46 46.84 53.06 50.00 57.58 58.90 48.42 54.70
jurBERT-large 45.16 47.30 69.23 45.79 47.96 54.17 68.69 67.12 56.84 55.81
Table 13: Detailed results for promotion exams.
Model Civil Penal Proc Civil Proc Penal Average
QBERT
RoBERT-small 29.41 49.10 37.25 33.33 37.27
RoBERT-base 27.45 45.09 27.45 25.49 31.37
RoBERT-large 29.41 37.25 37.25 35.29 34.80
jurBERT-base 41.18 33.33 33.33 31.37 34.80
jurBERT-large 31.37 21.57 33.33 39.22 31.37
CrossQBERT
RoBERT-small 21.57 49.02 43.14 37.25 37.75
RoBERT-base 31.37 27.45 27.45 25.49 27.94
RoBERT-large 41.18 29.41 29.41 31.37 32.84
jurBERT-base 25.49 35.29 35.29 42.14 34.55
jurBERT-large 35.29 41.18 43.14 27.45 36.77
ColBERT
RoBERT-small 23.53 33.33 35.29 27.45 29.90
RoBERT-base 33.33 29.41 27.45 21.57 27.94
RoBERT-large 21.57 43.14 50.98 31.37 36.77
jurBERT-base 31.37 33.33 43.14 41.18 37.26
jurBERT-large 45.10 41.18 29.41 31.37 36.77
LLM P1
FLAN-T5 XL 25.49 21.57 11.76 27.45 21.56
FLAN-T5 XXL 35.29 31.37 13.72 23.53 25.97
Mistral-Instruct v0.1 29.41 13.72 11.76 25.49 20.10
Mistral-Instruct v0.2 26.31 21.57 11.76 21.56 20.29
LLM P2
FLAN-T5 XL 29.41 21.57 13.72 27.45 23.04
FLAN-T5 XXL 31.37 25.49 15.69 25.49 24.51
Mistral-Instruct v0.1 33.33 21.57 9.80 25.49 22.55
Mistral-Instruct v0.2 39.21 19.61 11.76 21.57 23.04
LLM RAG
FLAN-T5 XL 31.37 17.65 21.57 23.53 23.53
FLAN-T5 XXL 39.22 27.45 21.57 27.45 28.92
Mistral-Instruct v0.1 11.76 7.84 5.88 15.69 10.29
Mistral-Instruct v0.2 43.14 23.53 21.57 27.45 28.92
LLM LFT
FLAN-T5 XL 35.29 45.10 41.18 41.18 40.69
FLAN-T5 XXL 37.25 43.14 35.29 29.41 36.27
Mistral-Instruct v0.1 19.61 29.41 29.41 27.45 26.47
Mistral-Instruct v0.2 17.65 17.65 17.65 23.53 19.12
GRAF(Ours)
RoBERT-small 43.14 52.94 45.10 45.10 46.57
RoBERT-base 60.78 58.82 41.18 50.98 52.94
RoBERT-large 52.94 62.75 54.90 56.86 56.86
jurBERT-base 47.06 52.94 50.98 45.10 49.02
jurBERT-large 54.90 54.90 52.94 49.01 52.94
Table 14: Detailed results for entrance exams.
Model Civil Penal Proc Civil Proc Penal Average
QBERT
RoBERT-small 37.90 32.26 33.06 32.26 33.87
RoBERT-base 26.61 27.42 29.84 36.29 30.04
RoBERT-large 34.68 32.26 40.32 33.87 35.28
jurBERT-base 33.87 23.58 33.06 33.06 30.89
jurBERT-large 40.32 31.71 32.26 30.65 33.74
CrossQBERT
RoBERT-small 34.68 31.97 39.52 34.68 35.21
RoBERT-base 30.65 38.52 29.84 30.65 32.42
RoBERT-large 41.94 36.07 33.87 34.68 36.64
jurBERT-base 22.58 34.43 29.84 34.68 30.28
jurBERT-large 36.29 27.05 31.45 30.65 31.36
ColBERT
RoBERT-small 29.03 33.06 28.23 27.42 29.44
RoBERT-base 38.71 27.42 34.68 29.84 32.66
RoBERT-large 36.29 32.26 36.29 39.52 36.09
jurBERT-base 23.39 29.03 31.45 35.48 29.84
jurBERT-large 29.03 33.87 35.48 33.87 33.06
LLM P1
FLAN-T5 XL 19.35 25.00 17.74 12.90 18.75
FLAN-T5 XXL 21.77 24.19 23.38 18.54 21.97
Mistral-Instruct v0.1 14.52 12.10 16.13 13.70 14.11
Mistral-Instruct v0.2 19.35 22.58 15.32 14.52 17.94
LLM P2
FLAN-T5 XL 17.74 22.58 19.35 13.71 18.35
FLAN-T5 XXL 18.55 24.19 23.39 13.71 19.96
Mistral-Instruct v0.1 14.52 18.55 16.13 12.90 15.53
Mistral-Instruct v0.2 15.32 24.19 16.13 14.52 17.54
LLM RAG
FLAN-T5 XL 21.77 22.58 20.16 16.94 20.36
FLAN-T5 XXL 20.97 26.61 33.87 20.97 25.61
Mistral-Instruct v0.1 5.65 11.29 12.90 5.65 8.87
Mistral-Instruct v0.2 26.61 29.03 20.16 19.35 23.79
LLM LFT
FLAN-T5 XL 31.45 32.26 29.03 36.29 32.26
FLAN-T5 XXL 27.42 28.23 33.06 22.58 27.82
Mistral-Instruct v0.1 27.42 31.45 21.77 29.03 27.42
Mistral-Instruct v0.2 13.71 14.52 24.19 23.39 18.95
GRAF(Ours)
RoBERT-small 46.77 39.52 43.90 43.44 43.41
RoBERT-base 48.39 49.19 52.85 46.72 49.29
RoBERT-large 50.00 47.58 49.59 55.73 50.73
jurBERT-base 51.61 55.65 51.22 59.02 54.38
jurBERT-large 51.61 61.29 56.10 57.38 56.60
Table 15: Detailed results for bar exams.
Appendix H Romanian Prompts
P1 - Entrance, Bar Răspunde la următoarea întrebare de legalitate din {tip drept}. Cel mult 2 răspunsuri sunt corecte. Dacă un singur răspuns este corect, vei răspunde doar cu litera răspunsului corect. Dacă 2 răspunsuri sunt corecte, vei răspunde doar cu literele răspunsurilor corecte: {tip drept} {întrebare} {variante de răspuns}
P1 - Promotion Răspunde la următoarea întrebare de legalitate din {tip drept}. Un singur răspuns este corect. Tu vei răspunde doar cu litera răspunsului corect: {tip drept} {întrebare} {variante de răspuns}
P2 - Entrance, Bar Răspunde la următoarea întrebare de legalitate din {tip drept}. Cel mult 2 răspunsuri sunt corecte. Dacă un singur răspuns este corect, vei răspunde doar cu litera răspunsului corect. Dacă 2 răspunsuri sunt corecte, vei răspunde doar cu literele răspunsurilor corecte Răspunde cu doar unul dintre simbolurile din lista [A, B, C, AB, AC, BC]: {tip drept} {întrebare} {variante de răspuns}
P2 - Promotion Răspunde la următoarea întrebare de legalitate din {tip drept} cu doar una dintre literele din lista [A, B, C]. Un singur răspuns este corect: {tip drept} {întrebare} {variante de răspuns}
[…]
Appendix I Translated Prompts
P1 - Entrance, Bar Answer the following legal question from {law type}. At most 2 answers are correct. If a single answer is correct, you will only answer with the letter of the correct answer. If 2 answers are correct, you will answer with only the letters of the correct answers: {law type} {question} {answer choices}
P1 - Promotion Answer the following legal question from {law type}. A single answer is correct. You will only answer with the letter of the correct answer: {law type} {question} {answer choices}
P2 - Entrance, Bar Answer the following question from {law type}. At most 2 answers are correct. If a single answer is correct, you will only a answer with the letter of the correct answer. If 2 answers are correct, you will answer with the letters of the correct letters. Answer with only one of the symbols from the list [A, B, C, AB, AC, BC]: {law type} {question} {answer choices}
P2 - Promotion Answer the following legal question from {law type} with only one of the letters from the list [A, B, C]. A single answer is correct: {law type} {question} {answer choices}
FLAN-T5 RAG - Entrance, Bar Answer the following legal question from {law type}. At most 2 answers are correct. If a single answer is correct, you will only answer with the letter of the correct answer. If 2 answers are correct, you will answer with only the letters of the correct answers If the information from the context is not in the question, then ignore the context and answer the question. Answer with only one of the symbols from the list [A, B, C, AB, AC, BC]. Context: {documets} Question: {question} {answer choices}
FLAN-T5 RAG - Promotion Answer the following legal question from {law type} with only one of the letters from the list [A, B, C]. If the information from the context is not in the question, then ignore the context and answer the question. A single answer is correct. Context: {documents} Question: {question} {answer choices}
Mistral RAG - Entrance, Bar Answer the following legal question from {law type}. At most 2 answers are correct. If a single answer is correct, you will only answer with the letter of the correct answer. If 2 answers are correct, you will answer with only the letters of the correct answers. If the information from the context is not in the question, then ignore the context and answer the question. Context: {documents} Question: {question} {answer choices}
Mistral RAG - Promotion Answer the following legal question from {law type}. You will only answer with the letter of the correct answer. If the information from the context is not in the question, then ignore the context and answer the question. A single answer is correct. Context: {documents} Question: {question} {answer choices}
Claim Graph Extraction - LLM Prompt Extract all entities and all relationships between entities from the legal text based on the example. At the end add STOP. You will answer with triplets of the form: (entity;relation;entity). The triplets are separated on lines. Each triplet relationship will be entered separately. Entities can be institutions, organizations, persons, functions, documents, courts and others. Text: (1) An assets investigation commission, hereinafter referred to as the investigation commission, shall operate in addition to each court of appeal, consisting of: a) 2 judges from the court of appeal, designated by its president, one of whom shall act as president, b) a prosecutor from the prosecutor’s office operating under the court of appeal, designated by the chief prosecutor of this prosecutor’s office. (2) The president and members of the investigation commission shall be designated for a period of 3 years. During the same period and by the same persons, 3 alternates will also be appointed, who will replace the holders in the event that they, for legal reasons, cannot take part in the work of the investigation commission. (3) The investigation commission has a secretary, appointed by the president of the court of appeal from among the clerks of this court. Entity;Relationship;Entity: (court of appeal;shall operated in addition to;assets investigation commission) (assets investigation commission;referred to as;investigation commission) (investigation commission;consisting of;2 judges) (2 judges;designated by;president of the court of appeal) (investigation commission;consisting of;prosecutor) (prosecutor;from;prosecutor’s office operating under the court of appeal) (prosecutor;designated by;chief prosecutor) (president of the investigation commission;designated for a period of;3 years) (members of the investigation commission;designated for a period of;3 years) (3 alternates;appointed by;the president of the court of appeal) (3 alternates;appointed by;the chief prosecutor) (3 alternates;designated for a period of;3 years) (3 alternates;will replace the holders if they cannot take part in the work of the investigation commission on;the heads) (investigation commission;has;a secretary) (a secretary;appointed by among the clerks of;the president of the court of appeal) STOP Text: {text} Entity;Relationship;Entity: